### 27 Mar 2026 v1

# Two-branch retention behavior in unsaturated fractured rock driven by fracture–matrix flow partitioning

Abstract
Upscaling unsaturated flow in fractured rock remains challenging because fractures and matrix often exhibit sharply contrasting hydraulic behaviors across saturation states. Here, we demonstrate that unsaturated flow undergoes a transition between matrix- and fracture-dominated regimes. Three-dimensional direct numerical simulations reveal that both relative permeability and capillary pressure curves display a robust two-branch structure. We analytically derive a generalized retention formulation that identifies a critical saturation marking the transition between the two distinct retention regimes and reproduces the two-branch behavior captured in the numerical simulations. An analytical expression for the critical pressure head is further derived to represent the limiting case of fully connected fracture networks, providing a physical explanation for the retention regime shift and showing good agreement with the numerical results for systems above the percolation threshold. Our results provide a mechanistic framework for understanding and upscaling unsaturated flow in fractured rock, with broad implications for hydrology and geophysics.
1 Introduction
Unsaturated flow in fractured rock plays a critical role in subsurface fluid transport, with important implications for groundwater management and nuclear waste disposal (Pruess & Tsang, 1990; Liu et al., 1998; Nimmo, 2012; Rutqvist & Tsang, 2012; Zhang, 2018; Wang et al., 2019; Liang et al., 2022). However, predicting unsaturated flow in such systems remains challenging, because fractured media consist of two distinct domains, i.e., the rock matrix and the fracture network, with strongly contrasting hydraulic and retention properties (Glass et al., 1995; Seol et al., 2003; Huang et al., 2005). Unlike saturated conditions, where fractures typically dominate fluid flow, under unsaturated conditions, the relative importance of matrix versus fracture flow remains unresolved.
From a theoretical perspective, the porous rock matrix is expected to play a dominant role under unsaturated conditions, because strong capillary forces promote imbibition into the fine pore space of the matrix, which can suppress rapid flow through fractures and lead to relatively diffuse infiltration (Wang & Narasimhan, 1985; Nitao & Buscheck, 1991). Based on this conceptual view, subsurface flow has been traditionally represented using continuum approaches that treat fractured rock as an equivalent porous medium (Peters & Klavetter, 1988). However, field observations have revealed a different behavior. Field experiments at sites such as Yucca Mountain (Nevada, United States) have shown that water can migrate rapidly through localized preferential pathways along fractures even under unsaturated conditions (Pruess, 1998; Pruess et al., 1999; Bodvarsson et al., 2003). Laboratory experiments have also illustrated that flow can proceed with minimal imbibition into the surrounding rock matrix, thereby limiting fracture–matrix hydraulic exchange and enabling fast fracture-dominated transport (Liu et al., 1998; Wood et al., 2002; Ji et al., 2004; Tokunaga et al., 2005; Zhou et al., 2022). Furthermore, the geometry and aperture distribution of fracture networks strongly influence recharge efficiency and the rate at which water percolates through the unsaturated zone (Liu et al., 2002; Cey et al., 2006; Zhu et al., 2021).
These concepts and findings expose a paradox: capillary forces favor matrix-dominated flow, yet fractures serve as preferential pathways for unsaturated flow. Here, we address this paradox by showing that flow in fractured rock exhibits a two-branch retention behavior, reflecting a transition between matrix-dominated and fracture-dominated regimes. Combining three-dimensional numerical simulations with analytical development, we derive a generalized retention formulation that captures this transition behavior and reveals the underlying physical mechanisms. These results provide a more complete and physically based picture of unsaturated flow in fractured rock.
2 Methodology
Unsaturated flow through fractured porous media can be described by the Richard’s equation (Bear, 1972). Under steady-state conditions, neglecting gravity and in the absence of sinks or sources, the governing equations for the matrix and fracture domains are respectively written as:
$\nabla\cdot\left(-\rho\frac{k_{m}k_{rm}}{\mu}\nabla p\right)=0$ (1)
$\nabla_{\tau}\cdot\left(-b\rho\frac{k_{f}k_{rf}}{\mu}\nabla_{\tau}p\right)=\rho\left(q^{+}+q^{-}\right)$ (2)
where $p$ is fluid pressure, $\rho$ is fluid density, $\mu$ is dynamic viscosity, $b$ is fracture aperture, $k$ is absolute permeability, $\nabla_{\tau}$ denotes the tangential gradient operator along the fracture plane, and $k_{r}$ is relative permeability, with the subscripts “m” and “f” indicating matrix and fracture properties, respectively. Fracture permeability $k_{f}$ is calculated based on the cubic law (Witherspoon et al., 1980), i.e., $k_{f}=b^{2}/12$. The term $\rho(q^{+}+q^{-})$ in Equation 2 represents fracture-matrix mass exchange, with the interfacial fluxes defined as $q^{+}=-(k_{m}k_{rm}/\mu)(\partial p^{+}/\partial n^{+})$ and $q^{-}=-(k_{m}k_{rm}/\mu)(\partial p^{-}/\partial n^{-})$, where $n^{+}$ and $n^{-}$ denote the outward normal directions on the respective fracture faces.
Unsaturated flow behavior is additionally controlled by constitutive relationships linking capillary pressure $p_{c}$, effective saturation $S_{e}$, and relative permeability $k_{r}$. The $p_{c}$$S_{e}$ relationship is commonly referred to as the retention curve. Here, we adopt the Brooks-Corey model (Brooks & Corey, 1966) with separate parameter sets specified for the fracture and matrix domains. Expressed in terms of pressure head $h$, the Brooks–Corey formulation is written as:
$S_{e}=\begin{cases}\dfrac{1}{\lvert\alpha h\rvert^{n}},&h<-\dfrac{1}{\alpha}\\[5.0pt] 1,&h\geq-\dfrac{1}{\alpha}\end{cases}$ (3)
$k_{r}=\begin{cases}S_{e}^{\frac{2}{n}+l+2},&h<-\dfrac{1}{\alpha}\\[5.0pt] 1,&h\geq-\dfrac{1}{\alpha}\end{cases}$ (4)
$p_{c}=\frac{\rho g}{\alpha}S_{e}^{-\frac{1}{n}}$ (5)
where $g$ is the gravitational acceleration, $\alpha$ is the inverse of the air entry pressure head, and $n$ and $l$ are constants. While several alternative retention models (van Genuchten, 1980; Rogers et al., 1983) exist, we adopt the Brooks-Corey formulation because it is well suited to rock materials (Assouline, 2005; Gershenzon et al., 2016). To demonstrate the generality of our framework, we additionally test the van Genuchten formulation, with results presented in Text S5 and Figure S12 in Supporting Information.
We construct a series of three-dimensional synthetic fracture networks within a cubic domain of side length $L$, where fracture radii $r$ follow a power-law distribution (Bonnet et al., 2001), with the density function given by:
$f(r,L)=\xi r^{-a},\quad r_{\min}\leq r<r_{\max}$ (6)
where $a$ is the power-law exponent, $\xi$ is a density term, and $r_{\min}$ and $r_{\max}$ are the minimum and maximum fracture radii. Fracture orientations and locations are assumed to be purely random, representing nominally isotropic conditions. Fracture networks are characterized by the fracture intensity $\gamma$ (defined as total fracture area per unit volume) and the percolation parameter $\chi$ (defined as normalized excluded volume), which are respectively related to the second and third moments of the density function as follows (Bour & Davy, 1998; de Dreuzy et al., 2000):
$\gamma(r^{\prime},L)=\int\frac{f(r,L)\pi r^{\prime 2}}{L^{3}}\,dr^{\prime}$ (7)
$\chi(r^{\prime},L)=\int\frac{f(r,L)\pi^{2}r^{\prime 3}}{L^{3}}\,dr^{\prime}$ (8)
where $r^{\prime}$ is the effective fracture radius corresponding to the portion of a fracture contained within the finite domain. The percolation parameter $\chi$ is a metric quantifying fracture network connectivity, where the system transitions from disconnected to connected states as $\chi$ goes beyond the percolation threshold $\chi_{c}$ having a value is around 0.7 to 2.8 (Balberg, 1985; de Dreuzy et al., 2000).
We consider five power-law exponents ($a=2.5$, 3, 3.5, 4, and 4.5) and five fracture intensities ($\gamma=2.5/L$, $5/L$, $10/L$, $20/L$, and $40/L$), resulting in 25 parametric combinations (Figure 1). Ten realizations are generated for each case. Fractures are assigned with a constant aperture of $L/1000$ and a porosity of unity (i.e., no infilling material). The surrounding matrix has a permeability of $10^{-18}\times L^{2}$ and a porosity of 0.01. Distinct Brooks-Corey parameters are assigned to fractures and matrix. For fractures, the inverse air-entry suction parameter $\alpha_{f}$ is related to fracture aperture, contact angle, and fluid properties through Jurin’s law (Ababou, 2018), expressed as follows:
$\alpha_{f}=\frac{b}{2\sigma_{w}\cos\theta}{\rho g}$ (9)
where $\sigma_{w}$ is the water-air surface tension and $\theta$ is the contact angle. The matrix pore radius is significantly smaller than the fracture aperture with $\alpha$ for matrix set to be $1/L$. The Brooks-Corey parameter $n$ was assigned to be 2 for the matrix and 1.5 for the fractures, consistent with reported lower values for fractures relative to the matrix (Li et al., 2014).
A series of steady-state simulations was conducted to explore different saturation states. To span the full range of effective saturation, the reference pressure head imposed at the inflow boundary is systematically varied using 126 logarithmically spaced values, ranging from 0.001L-0.001L to 2000L-2000L, with locally refined spacing over a selected interval to improve resolution near the transition. A uniform hydraulic gradient of 0.001-0.001 is imposed in the flow direction along the x-axis. The remaining four lateral boundaries were treated as impermeable. For each simulation, the resulting capillary pressure and relative permeability fields are extracted for subsequent analysis. Air pressure is assumed constant, such that negative water pressure corresponds directly to capillary pressure. The upscaled relative permeability is determined from the total volumetric outflow rate $Q$ at the outlet boundary, accounting for contributions from both fracture and matrix domains. The absolute permeability $k_{\mathrm{abs}}$ is obtained from a fully saturated simulation. By rearranging Equation 1, the relative permeability of the system can be obtained as:
$k_{r}=-\frac{Q\mu}{k_{\mathrm{abs}}L\Delta P}$ (10)
where $\Delta P$ is the pressure differential imposed across the system. All results and analyses in our study are presented in dimensionless form to ensure generality.
Figure 1: Generated three-dimensional fracture networks associated with different power-law exponents of fracture radii $a$ and fracture intensities $\gamma$. Ten realizations are generated for each case, with only one representative realization presented here for illustration. The percolation parameter $\chi$ is also computed to quantify fracture network connectivity.
3 Results
3.1 Emergent two-branch retention behavior
Figure 2 shows the upscaled relative permeability and capillary pressure curves for a representative fracture network case ($a=3$, $\gamma=20/L$; Figure 1), with the contributions from the fracture and matrix domains also shown. Both the relative permeability (Figure 2a) and normalized capillary pressure $p_{c}/\rho gL$ (Figure 2b) display a distinct two-branch structure, arising from the contrasting hydraulic behavior of the fracture and matrix domains across the saturation range. At low effective saturations, flow is dominated by matrix. As saturation increases, the system transitions to a regime dominated by fracture flow. The fracture contribution sharply increases at $S_{e}\approx 0.37$, referred to hereafter as the critical saturation $S_{c}$, which marks the transition from matrix-dominated to fracture-dominated regimes. This two-branch retention behavior is consistently observed across the broad range of fracture network scenarios examined in this study, including all stochastic realizations generated (Text S3 and Figures S8 and S9 in Supporting Information).
Figure 2: Numerical simulation results of upscaled (a) relative permeability $k_{r}$ and (b) normalized capillary pressure $p_{c}/\rho gL$ as functions of effective saturation $S_{e}$ in a fractured rock with the power-law exponent $a=3$ and fracture intensity $\gamma=20/L$. Matrix and fracture contributions are also separately shown to illustrate their relative roles. The black cross indicates the critical saturation $S_{c}$, marking the transition between matrix-dominated and fracture-dominated flow regimes.
We propose a two-branch Brooks-Corey model that captures the transition between matrix- and fracture-dominated flow regimes, given by:
$k_{r}(S_{e})=\begin{cases}k_{rc}\left(\dfrac{S_{e}}{S_{c}}\right)^{\frac{2}{n_{m}}+l_{m}+2},&S_{e}\leq S_{c}\\[7.5pt] k_{rc}+(1-k_{rc})\left(\dfrac{S_{e}-S_{c}}{1-S_{c}}\right)^{\frac{2}{n_{f}}+l_{f}+2},&S_{e}>S_{c}\end{cases}$ (11)
$p_{c}(S_{e})=\begin{cases}p_{cc}\left(\dfrac{S_{e}}{S_{c}}\right)^{-\frac{1}{n_{m}}},&S_{e}\leq S_{c}\\[7.5pt] p_{cc}-\left(\dfrac{\rho g}{\alpha_{f}}+p_{cc}\right)\left(\dfrac{S_{e}-S_{c}}{1-S_{c}}\right)^{-\frac{1}{n_{f}}},&S_{e}>S_{c}\end{cases}$ (12)
where $k_{rc}$ is the relative permeability at $S_{c}$ and $p_{cc}$ is the capillary pressure at $S_{c}$. This generalized formulation extends the classical Brooks-Corey relationship by introducing regime-dependent normalization and incorporating the critical saturation (associated with $k_{rc}$ and $p_{cc}$) separating the two regimes. If the effective saturation $S_{e}$ is below the critical saturation $S_{c}$, the system resides in the matrix-dominated regime, in which the constitutive relationships are expressed in terms of the normalized saturation $S_{e}/S_{c}$. For $S_{e}>S_{c}$, corresponding to the fracture-dominated regime, where the constitutive relationships are expressed in terms of the normalized critical saturation $(S_{e}-S_{c})/(1-S_{c})$. The constitutive parameters $n$ and $l$ are assigned with matrix properties for $S_{e}$ below $S_{c}$, and assigned with fracture properties for $S_{e}$ above $S_{c}$. In the limits $S_{c}\to 0$ and $S_{c}\to 1$, the generalized formulation reduces to the classical Brooks–Corey model for fracture-only and matrix-only flow, respectively. This two-branch analytical formulation closely matches the numerical simulation results and well captures the regime transition at $S_{c}$ (Figure 3).
Figure 3: Comparison of numerically simulated and analytically predicted relative permeability and capillary pressure curves. The scatter points represent simulation results of relative permeability $k_{r}$ and normalized capillary pressure $p_{c}/\rho gL$ as functions of effective saturation $S_{e}$, while the solid lines indicate the prediction by the two-branch analytical formulation.
3.2 Transition between matrix- and fracture-dominated regimes
As saturation increases, the matrix and fracture domains evolve differently due to the contrast in their air-entry pressures. The matrix reaches full saturation earlier and attains its maximum flow capacity with the relative permeability $k_{rm}$ reaching unity, while fractures remain partially saturated and their relative permeability $k_{rf}$ continues to increase. Once the matrix becomes fully saturated, the transition to fracture-dominated flow occurs at the point where fracture and matrix fluxes are equal, beyond which fractures dominate the flow.
By combining Equations 3 and 4, the relative permeability $k_{r}$ can be expressed as:
$k_{r}=\left(\frac{1}{|\alpha h|^{n}}\right)^{\frac{2}{n}+l+2}$ (13)
Substituting Equation 13 into Darcy’s law yields the volumetric flow rate as a function of pressure head $h$:
$Q=-\frac{kA}{\mu}\left(\frac{1}{|\alpha h|^{n}}\right)^{\frac{2}{n}+l+2}\frac{\Delta P}{L}$ (14)
Using Equation 14 and the condition $\int Q_{m}\,dL=\int Q_{f}\,dL$, the critical pressure head $h_{c}$ at this transition is derived as:
$h_{c}=\left[\frac{k_{f}\int A_{f}(L)\,dL}{k_{m}\int A_{m}(L)\,dL}\frac{\alpha_{m}^{2+n_{m}(l_{m}+2)}}{\alpha_{f}^{2+n_{f}(l_{f}+2)}}\right]^{\frac{1}{n_{f}(l_{f}+2)-n_{m}(l_{m}+2)}}$ (15)
where $A_{f}$ and $A_{m}$ denote the cross-sectional areas of the fracture and matrix domains, respectively, normal to the flow direction.
The total fracture cross-sectional area scales with the fracture intensity $\gamma$ and the aperture $b$, whereas the matrix occupies the remainder. Incorporating this geometric partitioning into Equation 15 yields the following analytical expression for the critical pressure head:
$h_{c}=\frac{1}{\alpha_{f}}\left[\frac{k_{f}b\gamma}{k_{m}(1-b\gamma)}\right]^{\eta}\approx\frac{1}{\alpha_{f}}\left(\frac{k_{f}b\gamma}{k_{m}}\right)^{\eta}$ (16)
where
$\eta=\frac{1}{2+n_{f}l_{f}+2n_{f}}$ (17)
Here, we use the approximation $(1-b\gamma)\approx 1$ given that $b\gamma\ll 1$. Combining Equations 9 and 16, and invoking the cubic law for fracture permeability, the critical pressure head $h_{c}$ can be expressed explicitly as a function of fracture aperture:
$h_{c}=\frac{b^{3\eta-1}2\sigma_{w}\cos\theta}{\rho g}\left[\frac{\gamma}{k_{m}(1-b\gamma)}\right]^{\eta}\approx\frac{b^{3\eta-1}2\sigma_{w}\cos\theta}{\rho g}\left(\frac{\gamma}{k_{m}}\right)^{\eta}$ (18)
The critical saturation $S_{c}$ can be further derived as (see Text S1 in Supporting Information for detailed derivations):
$S_{c}=\frac{\left|\frac{\alpha_{m}2\sigma_{w}\cos\theta}{b\rho g}\left(\frac{b^{3}\gamma}{12k_{m}(1-b\gamma)}\right)^{-\eta n_{m}}\right|(1-\gamma b)\phi_{m}+\left|\left(\frac{b^{3}\gamma}{12k_{m}(1-b\gamma)}\right)^{-\eta n_{f}}\right|\gamma b}{(1-\gamma b)\phi_{m}+\gamma b}$ (19)
where $\phi_{m}$ denotes the matrix porosity.
Figure 4 compares numerical results with the analytical solution (Equation 16) for the critical pressure head $h_{c}$. Here, we have rearranged Equation 16 to define a dimensionless critical pressure head $h_{c}\alpha_{f}/(b\gamma)^{\eta}$, which is plotted as a function of the percolation parameter $\chi$. For $\chi$ above the percolation threshold $\chi_{c}\approx 0.7-2.8$, the numerical results asymptotically approach the analytical solution, consistent with the assumption of fully connected fractures underlying Equation 16. Around and below the percolation threshold, systematic deviations emerge, because fracture connectivity becomes insufficient to sustain continuous flow pathways. Near the percolation threshold, stochastic variability in network connectivity also leads to coexistence of connected and disconnected realizations, explaining the scatter observed in the numerical results.
Figure 4: Numerical results (red dots) and analytical solution (black dashed line; Equation 16) for the dimensionless critical pressure head plotted against the percolation parameter $\chi$. The shaded gray area indicates the range of the percolation threshold $\chi_{c}\approx 0.7-2.8$.
4 Discussion
4.1 Mechanisms of retention regime transition in unsaturated fractured rock
The two-branch retention behavior revealed in this study reconciles the longstanding debate over the relative roles of matrix and fracture flow in unsaturated fractured rock. Our results show that both matrix and fractures play important roles, with their relative contributions strongly controlled by saturation. At low saturation, flow rates are small and capillary forces dominate; capillary imbibition draws water into the rock matrix, suppressing fracture flow and regulating air-water distribution (Demond & Roberts, 1991). As saturation increases, flow rates increase and viscous forces progressively outweigh capillary entry resistance (Xu et al., 1998); the matrix becomes fully saturated first with its relative permeability approaching unity, while fractures—owing to their weaker capillary effects—have not yet reached their maximum flow capacity, i.e., the relative permeability of fractures remains below unity (see Text S4 and Figures S10 and S11 in Supporting Information). As capillary pressure continues to decrease, the contribution of fracture flow to the total flow increases until it becomes comparable to, and eventually greater than, matrix flow. This competition between matrix and fracture flow gives rise to the two-branch retention behavior observed in our study, with the transition between the two regimes defined by a critical saturation $S_{c}$ at which fracture flow begins to exceed matrix flow. Despite previous experimental studies reporting distinct matrix- and fracture-dominated flow regimes (Nicholl et al., 1994; Tidwell et al., 1995; Liu & Bodvarsson, 2001; Glass et al., 2003; LaViolette et al., 2003; Rüdiger et al., 2024; Xiang et al., 2025), we here provide a quantitative framework that rationalizes these earlier observations and resolves the apparent paradox regarding the relative importance of matrix and fracture flow.
The retention regime transition discussed above occurs at a critical saturation $S_{c}$ associated with a critical pressure head $h_{c}$. Equation 16 shows that this critical saturation is primarily controlled by the permeability contrast between fractures and matrix. Because fracture permeability is governed by fracture aperture, variations in aperture strongly influence the critical head. Equation 16 further indicates that fracture retention parameters and inverse air-entry pressure directly affect the critical head, whereas matrix retention parameters do not play a role. Sensitivity analyses confirm this (Text S2 and Figures S1–S7 in Supporting Information). This is expected because the transition occurs after the matrix becomes fully saturated, when its relative permeability reaches unity and matrix retention parameters no longer play a role. It should be noted that this analytical prediction assumes well-connected fracture networks; in disconnected networks, matrix properties may also influence the transition.
To assess the generality of our framework, we have additionally tested the van Genuchten retention model (see Text S5 and Figure S12 in Supporting Information). A similar two-branch behavior is observed, with consistent values of $S_{c}$, $k_{rc}$ and $p_{cc}$ obtained, demonstrating that the framework is general and not restricted to the Brooks–Corey formulation.
4.2 Control of fracture network connectivity
The two-branch retention behavior persists across fracture networks with varying connectivity. Re-examination of Equation 16 shows that the critical transition is controlled primarily by the fracture intensity. The percolation parameter quantifying fracture network connectivity does not explicitly appear in the formulation, but it influences the validity of the analytical prediction. As shown in Figure 4, the analytical solution systematically underestimates the dimensionless critical pressure head for disconnected networks. This discrepancy arises from the assumption in Equation 16 that fractures form fully connected pathways for flow. While this assumption may hold for well-connected networks, where multiple flow paths distribute water flux, it breaks down near or below the percolation threshold (Davy et al., 2024). In disconnected networks, fractures alone cannot transmit flow through the system, and a substantial fraction of the flux must pass through matrix, thereby involving fracture–matrix exchange (Yoon et al., 2023), consistent with the limited increase in relative permeability between the critical saturation and full saturation (Text S2 and Figure S8 in Supporting Information). The fracture intensity used in Equation 16 therefore overestimates the hydraulically active fracture population in disconnected networks, leading to an underestimation of the dimensionless critical pressure head.
5 Conclusions
This study resolves the longstanding debate over the relative importance of matrix and fracture flow in unsaturated fractured rock. Using three-dimensional numerical simulations that explicitly represent interacting fracture–matrix systems, we show that unsaturated flow exhibits an emergent two-branch retention behavior reflecting a transition between matrix- and fracture-dominated regimes. A generalized analytical retention formulation is proposed, capturing the emergence of the two regimes and identifying a critical saturation that delineates them. The regime transition is revealed to be governed primarily by fracture properties, including aperture, air-entry pressure, and fracture intensity, whereas matrix constitutive parameters play a limited role once the matrix becomes fully saturated, particularly in well-connected networks. The two-branch retention behavior persists across fracture networks with varying connectivity, indicating that the regime transition arises fundamentally from fracture–matrix flow partitioning. These results provide a physically grounded basis for upscaling unsaturated flow in fractured rock by capturing the transition between matrix- and fracture-dominated regimes.
Acknowledgments
We are grateful for the financial support from the Mont Terri Consortium and Uppsala University. We also thank the Mont Terri PF-A experiment partners Swiss Federal Office of Topography (swisstopo), Swiss Federal Nuclear Safety Inspectorate (ENSI), German Federal Institute for Geosciences and Natural Resources (BGR), and Swiss Seismological Service (SED) for their support. Q.L. is grateful for support from the European Research Council (ERC) under the European Union’s Horizon Europe programme (ERC Consolidator Grant, grant no. 101232311) for the project “Unified framework for modelling progressive to catastrophic failure in fractured media (FORECAST)”.
Open Research
Derivations, benchmark simulations, sensitivity analyses for both single-fracture and fracture network cases, comparisons between numerical and analytical results, and additional evaluations of hydraulic parameter effects and retention models are provided in the Supporting Information.
Conflict of Interest Disclosure
The authors declare there are no conflicts of interest for this manuscript.